# More Modality, More AI: Exploring Design Opportunities of AI-Based Multi-modal Remote Monitoring Technologies for Early Detection of Mental Health Sequelae in Youth Concussion Patients

### v1

@@ 5.1. Procedure
We conducted user reviews with the same participants from our formative study. Due to the time constraints of concussion clinicians, each user review was limited to 10 minutes. We conducted a Zoom screen-sharing session to present our three modules to participants, introducing the data and functionalities included in each module. Based on each module, we then collected verbal feedback from participants regarding data type, visualization, interaction, usability, and potential use cases. Our research team documented their responses through note-taking. After the session, we collected participants’ feedback and applied axial coding to identify key themes. We then identified design modifications based on the themes. Next, we iterated on our initial system design following design modifications to come up with the final version of our system design. Our goal is to ensure that the final system design shows the potential to effectively support concussion clinicians in addressing the challenges of detecting mental health sequelae. The detailed interview protocol for our design interface evaluation is provided in Appendix B.
@@ 5.2. Findings
In this section, we presented the feedback of concussion clinicians on the initial system design, including the features they found helpful, their concerns and points of confusion, and suggested improvements to the system.
@@
5.2.1. AI Risk Score Prediction as Supportive Evidence for Clinicians in Patient Communication
@@
Concussion clinicians (P1, P5, P6) reported that if the AI risk score prediction module provides accurate data, it can be useful in assisting clinicians with treatment planning and making timely adjustments to clinical decisions, such as issuing referrals or a set of follow-up visits.
@@
”That [the AI risk score prediction] might be something helpful early on for them to say, Hey, this is somebody who probably gonna benefit from seeing another provider earlier rather than later.” (P5)
@@
Moreover, P1 mentioned that the AI Risk Score Prediction module is straightforward and has the potential to effectively support clinicians in communicating with patients and their parents when concerns arise about patients’ mental health sequelae during clinical visits. By directly presenting the risk scores of potential mental health outcomes, youth patients and their families gain a clearer understanding of both the current mental health state and the possibility of potential future mental health sequelae, particularly for youth patients and families who are hesitant to discuss mental health issues with concussion clinicians.
@@
”I think it’s pretty simple, and if I was able to show that to a family… maybe something as simple as this to show [the patient’] mom… ’I’m concerned about your son… There’s a 50% risk that may develop [mental health sequelae]. We should really be in touch with it.’ Yes, I think this is great.” (P1)
5.2.2. Explanation of AI Results and Data Sources as Supportive Information for Clinicians’ Understanding
@@
However, concussion clinicians expressed confusion regarding the data sources and the interpretation of results in certain modules. First, they raised concerns about the AI Risk Prediction module. When we presented this module, despite the interface providing key contributing data and a brief explanation of data sources, clinicians still found it difficult to understand the meaning of the risk score. They were also unsure whether the score was derived from other modules or from self-reported questionnaires administered to patients, such as the Generalized Anxiety Disorder 7-item Scale (GAD-7) and the Patient Health Questionnaire-9 (PHQ-9). Additionally, clinicians expressed confusion about the self-reported symptoms module. Specifically, P2 and P4 mentioned that they were unclear about how self-reported symptoms were collected.
@@
”So it’s just my confusion. So when it says self-reported symptoms that’s based on like a questionnaire that we’ve given them about their depression and anxiety level? Or this is what AI predicts that they have a 50 percenty chance of like developing depression based on that?” (P2)
5.2.3. CAs and Wearables as Additional Valuable Data Sources Outside the Clinic
@@
Concussion clinicians (P3, P6) expressed optimism about the potential of the self-reported symptoms module. They believed that patients might be more willing to share information with a voice assistant than directly with a clinician, making this module a valuable additional data source. One clinician mentioned that they would like to quickly review self-reported symptom data before each clinical visit to better understand youth patients’ current conditions. At the same time, P6 noted a potential challenge: if patients have a voice-based conversational agent at home, they might place it in a corner and forget about it. To address this, they suggested that CAs should initiate daily conversations proactively to collect patient information more effectively.
@@
”I guess from a perspective of maybe they would tell this device things that they wouldn’t tell me. Possibly I could see that.” (P6)
Among the three modules in our system design, concussion clinicians (P3, P4, P5) particularly appreciated the sleep and physical activity module. They highlighted that such data is currently missing from electronic health record (EHR) systems, yet wearable-generated data on sleep and physical activity could serve as key indicators for monitoring adolescent concussion patients’ mental health and recovery progress. P3 believed that this information could help monitor patients’ health on a weekly basis and assess whether they have followed clinical recommendations. However, P6 raised a concern that if a patient already owns a smartwatch, wearing a second device for physiological monitoring might be inconvenient. Additionally, having two devices could serve as a constant reminder of their injury, potentially reinforcing the perception that they have not yet fully recovered. Additionally, P3 suggested that it would be best if patients did not see the data collected by the smartwatch to avoid reminding them of their condition.
@@
”This one (the Sleep and Physical Activities panel) to me might be the most helpful [one]… I think that’s always hard [to get].” (P5)
”So I really like the sleep and physical activity, but I would suggest that the patients not see that data that it only goes to the physician. Because I don’t want it to cause them anxiety. So yeah, I really like that.” (P3)
5.2.4. Variations in Data Detail and Presentation Types for Clinicians’ Preferences
@@
Concussion clinicians exhibited varying preferences regarding the type and level of detail in the data presented within the sleep and physical activity module. Some clinicians preferred more detailed information than what was currently displayed. For instance, clinicians (P2, P3, P5) suggested including details related to napping under the sleep pattern section, such as whether the patient took naps and their duration. Additionally, clinicians (P1, P3) emphasized the importance of heart rate in assessing concussion recovery and monitoring mental health sequelae. Specifically, they expressed interest in viewing daily resting heart rate during sleep over the course of a week. They noted that during recovery, a patient’s heart rate tends to be lower, indicating a calm state, and it often correlates with obtaining eight hours of quality sleep. In contrast, if a patient experiences fragmented or restless sleep, their resting heart rate increases. By analyzing these patterns, clinicians could better assess whether a patient is experiencing anxiety.
@@
”If you (patients) are having a light, irritable kind of sleep, I would expect the resting heart rate to be in the seventies or eighties. [It means] maybe they (patients) are anxious. Maybe there’s something else going on. ” (P1)
@@
Moreover, in the self-report symptoms module, clinicians (P3, P5) suggested they want to include dietary information, such as appetite or nutrition, as they noted that when youth patients experience depression or high levels of anxiety, both the patients and their parents often report decreased appetite. Regarding the symptoms part of the module, instead of focusing solely on whether youth exhibit symptoms of mental health sequelae, concussion clinicians (P3, P5) emphasized that they are more concerned with the duration and severity of these symptoms. The severity and duration of symptoms have different impacts on clinicians’ treatment plans and decision-making, which potentially helps clinicians distinguish between general emotional issues and mental health sequelae following a concussion.
@@
”…persistent or non-improving symptoms. That’s what I look for.”(P5)
@@
Additionally, P3 expressed a new need for a feature that could indicate how adherence to recommendations impacts recovery rates, suggesting that such insights could enhance clinical decision-making.
”I don’t know if my patients are following my recommendations and knowing if they are and how that affects their recovery would be really interesting.” (P3)
5.2.5. Clinicians’ Concerns About Emergency Situation and Time Constraints
In the self-reported symptoms module, concussion clinicians placed significant emphasis on alerts related to suicidality. P1 emphasized the need for immediate intervention when suicidal tendencies emerge in youth concussion patients. However, they also highlight a critical gap in current clinical practice—the absence of established protocols for managing or intervening in such situations. When the user interface flags a potential suicide risk, clinicians report uncertainty about how to respond effectively due to a lack of concrete, actionable guidelines. Failure to intervene in a timely manner could result in missed opportunities for intervention and serious legal or ethical consequences. Additionally, even when clinicians recognize signs of suicidal ideation, managing emergency interventions without triggering resistance from parents or youth patients remains a highly sensitive and complex challenge. Furthermore, P4 reported that they face a dilemma: The youth generation experiences frequent mood swings and may sometimes exaggerate their feelings and symptoms. If a youth patient reports self-harm intentions that turn out to be overstated, clinicians worry about overreacting to false alarms. Striking the right balance between taking reports seriously and avoiding unnecessary escalations is a persistent concern in their practice. Clinicians (P2, P4) generally believe that parents should be brought into the loop when such situations arise.
@@ 5.2.5. Clinicians’ Concerns About Emergency Situation and Time Constraints
”So I think this is when you need to loop the paradigm and say, hey, your child has been identified through our software. You know, cause I think, especially since we’re dealing with minors, you’re gonna have to get the parents loop too.” (P4)
Moreover, clinicians (P2, P5) expressed concerns about their already heavy workloads, stating that they have little time to thoroughly review such detailed data. These clinicians noted that having dedicated staff to help filter and process data would greatly reduce their burden. Without such support, they often skim through relevant data quickly on the day of the appointment or shortly before, just to gain a basic understanding of the youth concussion patient’s condition.
@@
”I think that’s gonna that would be really tedious to try to monitor and create extra work for us.”(P5)
Regarding the newly introduced data and user interface modules, concussion clinicians (P1, P2) expressed concerns about the difficulty of finding time to process the data thoroughly. Clinicians already feel overwhelmed by the excessive features within current electronic health record (EHR) systems and their tasks. P1 highlighted that processing large amounts of data between tightly scheduled appointments—sometimes up to 20 patients per day—makes it nearly impossible to dedicate sufficient time and energy for in-depth analysis. Although monitoring systems can collect vast amounts of valuable data, clinicians often lack the capacity to fully review this information within their workflows. Given their heavy workload, the time clinicians can dedicate to each patient is extremely limited—sometimes as little as 15 minutes per appointment. As a result, concussion clinicians must rely on brief time windows, such as the moments before a scheduled meeting, to extract key information and quickly assess the condition of youth concussion patients.
”I would need to look at it on that day of the visit… Honestly, we are quite overwhelmed with the EHR. And all the different features and everything that we’re asked to do with it.”(P1)
6. Refined System Design
Based on the design interface evaluation, clinicians acknowledged the value of AI-RPM technologies which can potentially help them remotely collect and monitor patient data. In addition, clinicians found that the data, features, and presentation methods in our initial system interface could support their mental health-related decision-making. At the same time, they provided valuable feedback and suggestions based on the initial system design. In this section, we present design considerations derived from the design interface evaluation and use them to refine our initial system design. We then provide a refined system design in this section. The goal of presenting this refined system design is to visually illustrate the design considerations more effectively and provide insights for future studies.
6.1. Design Considerations
6.1.1. Explanation of Data Sources and Module Relationships
Through design interface evaluation, we found that while concussion clinicians considered the risk score useful for health decision-making and communication with patients and their families, they also raised concerns. Specifically, they were unsure how the score was generated and how it connected to other modules. To address this issue, we recommend using visualizations to illustrate the relationships between AI results and other data sources. This would enable clinicians to intuitively understand how various factors influence the risk score, potentially enhancing their trust in the system. One clinician explicitly mentioned wanting to see the relationship between clinical recommendations and patient recovery. So we suggest visualizing this by displaying the correlation between the risk score and exercise recommendation compliance.
6.1.2. Personalized Time Selection
One clinician found the AI-generated risk score unhelpful because the initial UI displayed a one-year risk score, whereas she typically treats youth patients for only four weeks. If a patient does not recover within this period, they are referred for further evaluation. Therefore, we believe clinicians would benefit from viewing AI model risk scores over different time spans to meet their varying needs and align with their workflow. To accommodate different clinical needs, we propose adding a time range selector to the module. This feature would allow clinicians to customize the risk score display based on their specific decision-making timeframes, offering personalized support for different clinical workflows.
Furthermore, for the sleep and activities module in Section 5.2.3, a clinician expressed a preference for viewing data on a weekly basis. To align with this preference, we recommend that the module default to displaying seven days’ worth of data while also providing the option to explore detailed data for specific dates. This would allow clinicians to closely examine anomalies, particularly when irregular patterns appear on certain days.
6.1.3. Providing Information on Other Stakeholders to Support Collaboration
Concussion clinicians noted that in the self-reported symptoms module, the most critical yet challenging situation for clinicians is detecting suicidality in patients. However, they expressed uncertainty about how to remotely respond if a patient exhibits self-harm tendencies. Additionally, clinicians raised concerns about the authenticity of such reports, as youth patients may sometimes exaggerate their symptoms. Clinicians emphasized that if an emergency occurs, parents should be involved, as they are considered key stakeholders in managing such situations. In response to this concern, we propose integrating an emergency contact module within this UI module, which allows clinicians to contact parents when necessary. Since parents may have quicker access to their children and be able to take immediate action, this feature could enhance response efficiency. Additionally, clinicians expressed a need to track the number of days symptoms persist, as this directly reflects a patient’s recovery progress and informs mental health-related decisions such as giving a referral. We recommend that when CAs engage in conversations with patients and their parents, they also document their observations on the duration of mental health symptoms. Then from the design side, we propose adding a data point in the module to display symptom duration.
In addition to monitoring the patient’s emotional state, clinicians often ask parents about behavioral changes. To enhance data comprehensiveness and objectivity, we suggest incorporating parent-VA (virtual assistant) conversation data into this module, offering an additional perspective on the patient’s condition. Finally, to reduce confusion and improve the module’s usability for clinicians, we need to clearly define how the data is collected. Ensuring transparency in data sources will help clinicians trust and effectively utilize this module in their practice.
6.1.4. Balancing Additional Data Presentation and Cognitive Load
In the initial system design, certain key data types that are highly important to clinicians were not covered. As section 5.2.4 mentioned, concussion clinicians emphasized the need for heart rate during sleep, nap times, wake-up time, and sleep time rather than just total sleep hours. To address this, we recommend displaying these additional data points within the module. However, given clinicians’ concerns about data overload—especially as they noted that existing electronic health record (EHR) systems have already increased their workload—we propose a customizable data selection feature. Within the system interface, clinicians can personalize their data view based on their specific needs and workload at different times, rather than being presented with all data at once. Additionally, we believe it is essential to prioritize the most relevant information in the default display. Specifically, the most critical data will be shown by default, while supplementary explanatory details will be secondary and only accessible when needed. This design approach aims to reduce cognitive load, ensuring that clinicians can quickly access key insights without being overwhelmed by excessive information. By minimizing cognitive burden, this method enables more efficient and personalized analysis.
6.1.5. Integrating Related Data into a Unified System Interface
In our formative study, clinicians mentioned using GAD-7 and PHQ-9 to monitor patients’ anxiety and depression. Additionally, as Section 5.2.2 mentioned, clinicians noted that they use these questionnaires to patients again during the design interface evaluation, which makes us realize that we could integrate all the information clinicians want to see into the same system interface. Based on this, we propose adding a questionnaire results module to the system interface, displaying patients’ completed questionnaires before their clinical visits. This integration would allow clinicians to access all relevant data in a single interface before a consultation, saving time, improving data review efficiency, and enabling a faster and more evidence-based assessment of patients’ mental health status. Importantly, these questionnaire results differ from the anxiety and depression statuses identified in the self-reported symptoms module. They provide a more structured and standardized assessment. By combining subjective self-reported symptoms with objective questionnaire scores, clinicians can gain a more comprehensive understanding of patients’ mental health conditions, which would support more confident mental health-related decision-making, even when patients are outside the clinical setting.
6.2. Visualization of Refined System Design
In this version, we considered the possibility of seamlessly integrating the system design with existing electronic health record (EHR) systems in the future. Therefore, we referenced the design of current EHR systems to ensure consistency. To align with the visual and interaction patterns of EHR systems, we incorporated a patient list on the left side of the interface. A red dot next to a patient’s name indicates a potential risk of mental health sequelae, making it easier for clinicians to identify at-risk patients at a glance. The final system design consists of five modules, which aim to enhance system usability and support clinical decision-making. In the following sections, we will explain the data, functionality, potential interactions, and purpose of each module in detail. The final system design interface is shown in Fig. 2.
Figure 2. The final system design is based on feedback from concussion clinicians gathered through design interface evaluation: (A) patient basic information, (B) patient sleep and physical activities, (c) patient self-report symptoms, (D) patient questionnaire result, and (E) AI risk score prediction for mental health sequelae.
6.2.1. The Basic Patient Information Module
To better integrate seamlessly with the existing EHR system, we have added a basic patient information section in the upper-left module of our system (Figure 2A). This section includes key details such as the date of injury, the number of days since the concussion occurred, the youth patient’s mental health history, and medication usage. These pieces of information provide clinicians with essential insights into the patient’s condition, including the time elapsed since the injury and their mental health background. In particular, the latter two details—mental health history and psychiatric medication usage—are important for concussion clinicians when detecting patients’ mental health sequelae.
6.2.2. The Sleep and Physical Activities Module

### v2

@@ 5.1. Procedure
We conducted an evaluation study with the same six clinicians from our formative study since they are the target users of the AI-RPM system. Due to the significant time constraints faced by concussion clinicians, each evaluation session was limited to 10 minutes with each clinician. We conducted a Zoom screen-sharing session to present our preliminary designs to each clinician and collected verbal feedback regarding data type, visualization, interaction, and usability. Meanwhile, our research team documented their responses through audio recording and note-taking. After the session, two researchers applied axial coding to identify codes and key themes. Afterward, we iterated on our preliminary designs based on our findings from this design interface evaluation. The detailed interview protocol for our evaluation study is provided in Appendix B.
@@ 5.2. Findings
In this section, we present five key findings derived from concussion clinicians’ feedback during the evaluation study.
@@
5.2.1. Improving Patient-Clinician Communication With AI Risk Score Prediction
@@
Concussion clinicians (P1, P5, P6) reported that if the AI risk score prediction module provides accurate data, it can be very useful for timely clinical adjustments, such as scheduling follow-up visits or providing referrals.
@@
“That (the AI risk score prediction) might be something helpful early on for them (concussion clinicians) to say, hey, this is somebody who probably gonna benefit from seeing another provider earlier rather than later.” (P5)
@@
Moreover, P1 mentioned that the AI risk score prediction module is not only easy to understand but also has the potential to serve as an evidence-based tool to effectively support clinicians in communicating with patients and their parents about mental health conditions during clinical visits. By directly presenting the risk scores of the likelihood of developing mental health sequelae, youth patients and their families may gain a clear understanding of the severity of patients’ mental health conditions and be willing to talk about it, particularly for youth patients and families who are hesitant to discuss mental health issues with concussion clinicians.
@@
“I think it’s pretty simple, and if I was able to show that to a family… maybe something as simple as this to show [the patient’s] mom… ’I’m concerned about your son… there’s a 50 percent risk that [your son] may develop [mental health sequelae]. We should really be in touch with it.’ Yes, I think this is great.” (P1)
However, P3 found the AI-generated risk score unhelpful because it covered one year, while she only treated youth concussion patients for four weeks. If patients did not recover within this period, they would be referred for further treatment.
5.2.2. Supporting Clinicians’ Understanding with AI-Generated Results and Data Source Transparency
@@
Concussion clinicians expressed confusion regarding the data sources and AI results in both the AI risk score prediction modules and self-reported symptoms modules. When we presented the AI risk score prediction module, P6 sought for a clear explanation of the meaning behind the AI risk score despite the interface already providing key data that contribute to the risk score as well as a brief explanation of the score. Additionally, P2 and P4 mentioned that they were unclear about how patient’s mental health symptoms were collected in the self-reported symptoms module. Specifically, P2 was unsure whether the symptoms (e.g., suicidality and depression), were from AI risk score prediction modules or from questionnaires (e.g., GAD-7 and PHQ-9) that clinicians gave to their patients for self-assessment. However, the symptoms are collected by LLM-powered CAs through check-in conversations with youth patients.
@@
“So it’s just my confusion. So when it says self-reported symptoms that’s based on like a questionnaire that we’ve given them about their depression and anxiety level? Or this is what AI predicts that they have a 50 percent chance of like developing depression based on that?” (P2)
5.2.3. Enriching Clinicians’ Insights with Data from LLM-based CAs and Wearables
@@
Clinicians reported that the data collected from CAs and wearables could be valuable in their workflow. Specifically, P6 mentioned that he would have a quick review of data on the self-reported symptoms module collected by CAs before each clinical visit to better understand youth patients’ current health conditions. Moreover, concussion clinicians (P3, P6) believed that patients might be more willing to share information with a conversational agent than directly with a clinician, which makes the data collected by CAs a valuable additional data source. However, P3 suggested that rather than using CAs to ask patients about their mental health conditions daily, it would be more appropriate to collect the data weekly. This suggestion helps avoid constantly reminding patients of their injuries.
@@
“I guess from a perspective of maybe they (patients) would tell this device things that they wouldn’t tell me. Possibly I could see that.” (P6)
Moreover, concussion clinicians (P3, P4, P5) particularly appreciated the data in the sleep and physical activities module collected by wearables. They highlighted that data such as sleep hours is currently missing from EHR systems, yet wearable-collected data could serve as key indicators for mental health sequelae and concussion recovery progress. Additionally, P3 believed that clinicians could use the sleep and activity data to assess whether patients have followed sleep and exercise recommendations, which can be useful for addressing one of their main challenges.
@@
“This one (the Sleep and Physical Activities panel) to me might be the most helpful [one]… I think that’s always hard [to get].” (P5)
5.2.4. Diverse Clinician Preferences in Data Details and Presentation Styles
@@
Concussion clinicians showed varying preferences regarding the type and level of detailed data presented within the sleep and physical activity module. Clinicians (P2, P3, P5) suggested incorporating napping-related data into the module, including whether the patient took naps and their duration. Additionally, clinicians (P1, P3) emphasized the importance of heart rate in assessing concussion recovery and monitoring mental health conditions. Specifically, they expressed interest in viewing daily resting heart rate during sleep over the course of a week. P1 noted that a heart rate in the fifties to sixties indicates a calm state in patients. A calm state is often associated with obtaining eight hours of quality sleep. In contrast, if a patient experiences fragmented or restless sleep, their resting heart rate increases. By analyzing physiological data, clinicians can assess if symptoms are linked to mental health or other non-concussion issues.
@@
“If someone’s recovering, I would expect it (heart rate) to be low, [meaning] they’re calm……If you (patients) are having a light, irritable kind of sleep, I would expect the resting heart rate to be in the seventies or eighties. [It means] maybe they (patients) are anxious. Maybe there’s something else going on.” (P1)
@@
Moreover, in the self-report symptoms module, clinicians (P3, P5) suggested they want to include dietary information, such as appetite or nutrition, as they noted that when youth patients experience depression or high levels of anxiety, both the patients and their parents often report decreased appetite. Importantly, instead of focusing solely on whether youth patients have mental health symptoms, concussion clinicians (P3, P5) were more concerned with the duration and severity of these symptoms. Because the severity and duration of symptoms have different impacts on clinicians’ decision-making, such as treatment plans.
@@
“I don’t think it has to be high level [of mental health symptoms]…… persistent or non-improving symptoms. That’s what I look for.” (P5)
@@
Additionally, P3 expressed a new need for a feature that could indicate how adherence to sleep and exercise recommendations impacts patients’ recovery, suggesting that such insights could enhance clinical decision-making.
5.2.5. Concerns of Clinicians in Emergency Situations and Time Constraints
In the self-reported symptoms module, concussion clinicians placed significant emphasis on alerts related to suicidal tendencies. P1 emphasized the need for immediate intervention when suicidal tendencies emerge in youth concussion patients. However, clinicians expressed several challenges. Firstly, there is an absence of established protocols for managing such emergency situations. Furthermore, P4 mentioned that the youth generation experiences frequent mood swings and may sometimes exaggerate their feelings, which could lead to false alerts. However, failure to intervene in patients who have suicidal tendencies in a timely manner could result in a life-or-death consequence. To address this, clinicians (P2, P4) believed that youth patients’ parents should be brought into the loop and have a partnership with parents when emergency situations arise.
@@ 5.2.5. Concerns of Clinicians in Emergency Situations and Time Constraints
“So I think this is when you need to loop the paradigm and say, hey, your child has been identified through our software. You know, cause I think, especially since we’re dealing with minors, you’re gonna have to get the parents loop too.” (P4)
Moreover, clinicians (P2, P5) expressed that they have limited time to thoroughly review such detailed data due to heavy workloads. Specifically, P1 highlighted that processing large amounts of data between tightly scheduled appointments—sometimes up to 20 patients per day—makes it nearly impossible to dedicate sufficient time and energy for in-depth analysis. If they have the AI-RPM system, clinicians noted that they might skim through the most relevant data quickly on the day of the appointment or shortly before to gain a basic understanding of their patient’s condition.
@@
“I think that’s gonna that would be really tedious to try to monitor and create extra work for us.” (P5)
6. Refined AI-RPM System Design
Based on the findings in the design interface evaluation session, clinicians acknowledged the value of the system with AI-RPM technologies that can potentially help them remotely monitor patients’ mental health symptoms and support them to make decisions in their workflow. At the same time, they provided feedback and suggestions based on the preliminary designs. Based on their feedback, we revised our preliminary design and proposed a complete design of clinician-facing AI-RPM system design (Fig. 1).
6.1. Reducing Cognitive Burden Through EHR Integration
To reduce the cognitive burden placed on clinicians by the new system and ensure the system can be seamlessly integrated into clinicians’ workflow, we referenced the style of current EHR systems to ensure consistency. We integrated a patient list (Fig. 1 A), which already exists in the EHR system, into the left side of the interface. However, we added a red dot next to a patient’s name to capture clinicians’ attention if AI-RPM technologies detect abnormalities in the patient’s mental health conditions. Moreover, we added a basic patient information section, which exists in the current EHR system, in the upper-left module (Fig. 1 B). However, the information here is tailored to concussion clinicians’ needs. These pieces of information provide clinicians with essential insights into a concussion patient’s condition. In particular, the Individual Psychiatric History and Mental Health Prescription are important for concussion clinicians when detecting youth patients’ mental health sequelae and making decisions.
6.2. The Sleep and Physical Activities Module
Based on findings from our evaluation study, concussion clinicians affirmed the value of the data provided in the Sleep and Physical Activities module (Fig. 2 C) in the preliminary design. However, several clinicians requested more information about data sources. In response, we designed a question mark next to the module title (Fig. 1 C). Concussion clinicians can hover over the question mark to view the data source. Moreover, because clinicians typically prefer a weekly overview of patient data, and some of them also want the ability to see daily metrics (e.g., heart rate), we provided the option to view data from both the last week and the previous 24 hours. With this option, concussion clinicians can easily track changes across different time intervals.